OpenStudy (ashia):

Solve the Proportion for x: 2/3 = x/30

OpenStudy (shamil98):

Cross multiply the terms.

OpenStudy (anonymous):

or multiply 30 to each side

OpenStudy (shamil98):

\[\huge \frac{ 2 }{ 3 } = \frac{ x }{ 30 } ---> 3x = 60\] now just divide 3 to both sides to get x.

OpenStudy (ashia):

thank you guys. i finally understand :)

OpenStudy (ashia):

will it be 20?

OpenStudy (shamil98):

yes

OpenStudy (ashia):

yes. thank you !!

OpenStudy (ashia):

how about this one ? i got 12a + 2 / 3a

OpenStudy (anonymous):

factor the first denominator, and find the LCD for them, then add them

OpenStudy (anonymous):

still need help with it or you're okay

OpenStudy (shamil98):

\[\huge \frac{ x+1 }{ x^2 + 2x + 1 } + \frac{ 1 }{ x+1 }\] Simplify the first fraction: \[\huge \frac{ \cancel{x+1}(1) }{ \cancel{(x+1)}(x+1) } + \frac{ 1 }{ (x+1) }\] \[\huge \frac{ 1 }{ x+1 } +\frac{ 1 }{ x+1 }\] \[\huge \frac{ 2 }{ x+1 }\]

OpenStudy (ashia):

thanks guys. im all set. thank you for your help:)